# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v2

@@ 4 Experiments
In Sec. 4.2, we validate DyGMamba’s ability in CTDG representation learning by comparing it with baseline methods on dynamic link prediction [^2] To supplement, we also validate on the dynamic node classification task. Since current mainstream datasets of this task requires no long-term temporal reasoning, we put the discussion in App. F. We do further analysis to show the effectiveness of model components, in particular the time-level SSM block. In Sec. 4.3, we further show DyGMamba’s efficiency. We also show that it achieves much stronger scalability in modeling long-term temporal information compared with DyGFormer.
@@ 4.1 Experimental Setting
We use the implementations and the best hyperparameters provided by Yu et al. (2023) for all baseline models except CTAN. For CTAN, we use its official implementation, fixing the number of layers to 5. All models are trained with a batch size of 200 for fair efficiency analysis. For DyGMamba, we report the number of sampled one-hop temporal neighbors $\rho$ and the patch size $p$ here. On Wikipedia, Social Evo., and UCI, $\rho\ \&\ p=$ 32 & 1. On Reddit, $\rho\ \&\ p=$ 64 & 2. On MOOC, $\rho\ \&\ p$ = 128 & 4. On Enron, $\rho\ \&\ p=$256 & 8. On LastFM, $\rho\ \&\ p=$ 512 & 16. Note that to fairly compare DyGMamba’s efficiency with DyGFormer, we keep the sequence length $\rho/p$ input into the SSM as same as the length input into Transformer in Yu et al. (2023), i.e., $\rho/p=$ 32. All experiments are implemented with PyTorch (Paszke et al., 2019) on a server equipped with an AMD EPYC 7513 32-Core Processor and a single NVIDIA A40 with 45GB memory. We run each experiment for five times with five random seeds and report the mean results together with error bars. Further implementation details including complete hyperparamter configurations are presented in App. C. We employ two evaluation settings following previous works: the transductive and inductive settings. As suggested in (Poursafaei et al., 2022), we do link prediction evaluation using three negative sampling strategies (NSSs): random, historical and inductive. Historical NSS is only considered under the transductive setting. See App. D for detailed explanations. We employ two metrics, i.e., average precision (AP) and area under the receiver operating characteristic curve (AUC-ROC)
@@ 4.2 Performance Analysis
Comparative Study on Benchmark Datasets.
@@
We conduct two ablation studies to study the effectiveness of model components. In study A, we make a model variant (Variant A) by removing the time-level SSM block and restrain our model from learning temporal patterns (information selection is substituted by mean pooling over the output of Eq. 5). In study B, we make a model variant (Variant B) by removing the Mamba SSM layers (Eq. 4) in the node-level SSM block. We find that: (1) From Table 3 and 4, we observe that under any NSS setting, Variant A is constantly beaten by the complete model, proving the effectiveness of learning edge-specific temporal patterns; (2) We also witness a performance drop from Table 3 and 4 if we switch DyGMamba to Variant B. This indicates the importance of encoding the one-hop temporal neighbors with SSM layers for capturing graph dynamics.
@@
A Closer Look into Temporal Pattern Modeling.
@@
We observe from ablation studies that time-level SSM block contributes to better model performance on real-world datasets. To better quantify the benefits brought by temporal pattern modeling, we construct three synthetic datasets, i.e., S1, S2 and S3, that follow different patterns and compare our model with DyGFormer, CTAN as well as Variant A on them. Each synthetic dataset contains 7 nodes, where the interactions of each pair of two nodes follow a certain pattern along time. And for each node, we generate interactions with all the other nodes. Assume we have a pair of node $u$ and $v$ and they have interactions at $\{t_{i}\}_{i=0}^{N}$, in S1, the time intervals between neighboring interactions $\{t_{1}-t_{0},...,t_{N}-t_{N-1}\}$ follow an increasing trend with a constant velocity of 0.05, i.e., $(t_{i+2}-t_{i+1})-(t_{i+1}-t_{i})=$ 0.05 . In S2, we set the time intervals to a decreasing trend with the same velocity, i.e., $(t_{i+1}-t_{i})-(t_{i+2}-t_{i+1})=$ 0.05. And in S3, we modify S1 by repeating several periods of increasing patterns taken from S1 to form a periodic dataset. In this way, we have three datasets demonstrating diverse temporal patterns: increasing/decreasing/periodic time intervals between neighboring interactions. Details of dataset construction and statistics are provided in App. A.2. From Table 5, we observe that DyGMamba greatly outperforms DyGFormer and CTAN. More importantly, Variant A shows similar performance to DyGFormer, meaning that our time-level SSM block is able to capture temporal patterns and modeling such patterns is important in CTDG reasoning. For more implementation details on synthetic datasets, please refer to App. C.2.
@@
Datasets DyGFormer CTAN Variant A DyGMamba
S1 55.19 $\pm$ 0.98 51.25 $\pm$ 2.11 53.72 $\pm$ 0.04 81.58 $\pm$ 1.31
S2 57.80 $\pm$ 4.61 51.17 $\pm$ 0.93 60.16 $\pm$ 2.20 85.36 $\pm$ 2.55
S3 79.20 $\pm$ 0.60 51.46 $\pm$ 0.19 76.36 $\pm$ 0.45 86.59 $\pm$ 0.09
@@
Datasets DyGFormer CTAN Variant A DyGMamba
S1 56.27 $\pm$ 0.54 51.25 $\pm$ 2.38 53.16 $\pm$ 0.39 86.61 $\pm$ 1.30
S2 59.06 $\pm$ 6.07 51.46 $\pm$ 0.91 62.50 $\pm$ 2.28 89.94 $\pm$ 2.70
S3 82.89 $\pm$ 1.34 52.12 $\pm$ 0.44 81.78 $\pm$ 0.40 91.72 $\pm$ 0.11
@@ 4.3 Efficiency Analysis
Figure 2: Efficiency comparison on four datasets among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory. The performance metric here is AP of transductive link prediction under random NSS. The greener, the better performance/efficiency. In contrast to other methods, DyGMamba consistently shows strong overall capability across different datasets. More explanations in Sec. 4.3.
@@
Model Size, Per Epoch Training Time and GPU Memory.
@@
Fig. 2 compares DyGMamba with five baselines in terms of number of parameters (model size), per epoch training time and GPU memory consumption during training [^3] The baselines not included here are either extremely inefficient (e.g., CAWN) or inferior in performance (e.g., DyRep). Complete statistics of all baseline models presented in App. G. We find that: (1) DyGMamba uses very few parameters while maintaining the best performance, showing a strong parameter efficiency. Only CTAN constantly uses fewer parameters than DyGMamba, however, its performance is much worse; (2) DyGMamba is much more efficient than DyGFormer with the same length of input sequence ($\rho/p=$32); (3) Although DyGMamba generally consumes more GPU memory and per epoch training time compared with most baselines, the gap of consumption is not very large in most cases. To model more temporal neighbors for long-range temporal dependent datasets, DyGMamba naturally requires more computational resources, thus enlarging the consumption gap. DyGFormer shows the same trend as DyGMamba since it also captures long-term temporal dependencies by considering more temporal neighbors; (4) CTAN requires very few computational resources. However, on long-range temporal dependent datasets, it is beaten by DyGFormer and DyGMamba with a great margin. Besides, CTAN is also hard to converge. Although it requires very little per epoch training time, it requires much more epochs to reach the best performance, leading to a long total training time. See App. G.2 and G.3 for details.
Impact of Patch Size on Scalability and Performance.
Patching treats $p$ temporal neighbors as one patch and thus decreases the sequence length by $p$ times. This is very helpful in cutting the consumption of GPU memory and training/evaluation time. However, patching introduces excessive parameters because it is done through $f_{N}$, $f_{E}$, $f_{T}$ and $f_{F}$ (Eq. 3) whose sizes increase as the patch size grows. Fig. 3b shows the numbers of parameters of DyGFormer and DyGmamba with different patch sizes on a long-term temporal dependent dataset Enron. We find that patching greatly affects model sizes. To further study how patching affects DyGMamba, we decrease the patch size gradually from 8 to 1 and track DyGMamba’s performance (Fig. 3a) as well as efficiency (Fig. 3b to 3d) on Enron. Meanwhile, we also keep track on DyGFormer under the same patch size for comparison. We have several findings: (1) Whatever the patch size is, DyGMamba always consumes fewer parameters, less GPU memory and per epoch training time, showing its high efficiency; (2) While both models require increasing computational budgets as the patch size decreases, the speed of increase is much lower for DyGMamba, demonstrating its strong scalability in modeling longer sequences; (3) Different trends in performance change are observed between two models. While DyGFormer performs worse, DyGMamba can benefit from a smaller patch size, indicating its strong ability to capture nuanced temporal details even if the sequence becomes much longer. Note that the models use fewer parameters under smaller patch size. This also shows that DyGMamba can achieve much stronger parameter efficiency by reducing patch sizes. To further study the reason for finding (3), we plot the performance of Variant A (from ablation study A) under different patch sizes in Fig. 3a. We find that Variant A’s performance degrades when sequence length is more than $64$. This means that dynamic information selection based on edge-specific temporal patterns is essential for DyGMamba to optimally process long sequences.
@@
Scalability in Modeling Increasing Number of Temporal Neighbors.
Sequence length is decided by the sampled temporal neighbors $\rho$ and the patch size $p$. If we want to model a huge
@@
number of temporal neighbors, e.g., $\rho=$4096, keeping the sequence length unchanged as 32 means we need to set $p$ to 128. This will greatly increase model parameters and cause burden in parameter optimization. Besides, as shown in Fig. 3a, bigger $p$ does not necessarily lead to better performance. Therefore, it is also important to see if a model is scalable to $\rho$, with a fixed $p$. In Fig. 4, we show the consumed GPU memory of DyGMamba and DyGFormer with $\rho$ varying from 64, 128, 256, 512, 1024, 2048, 4096 and 8192 on Enron, under a fixed patch size $p=$8. We find that DyGMamba shows superior scalability over DyGFormer. While DyGFormer can only process 2048 nodes, DyGMamba can deal with more than 8192 (at least 4 times) on a single 45GB NVIDIA A40 GPU. This implies our model’s potential to process a huge number of temporal neighbors with a limited GPU memory budget.
DyGMamba vs. DyGFormer on Complexity.
In historical interaction modeling, DyGMamba holds a computational complexity linear to $\rho/p$, while DyGFormer’s complexity is quadratic to $\rho/p$. As a result, as the sequence length grows (either $\rho$ increases or $p$ decreases), DyGFormer is less scalable compared with DyGMamba.

### v3

@@ 4 Experiments
In Sec. 4.2.1, we validate DyGMamba’s ability in CTDG representation learning by comparing it with baseline methods on dynamic link prediction [^3] To supplement, we also validate on the dynamic node classification task. Since current mainstream datasets of this task requires no long-term temporal reasoning, we put the discussion in App. F. Additionally, we also benchmark DyGMamba on DTDGs in App. L. This serves as supplementary experiment and does not directly connect to our main focus.. We show the effectiveness of model components by conducting ablation studies (Sec. 4.2.2) and analysis on synthetic datasets (Sec. 4.2.3). In Sec. 4.3.1, we show DyGMamba’s efficiency against various baselines. We also show that it achieves much stronger scalability in modeling long-term temporal information compared with the current state-of-the-art DyGFormer (Sec. 4.3.2 and Sec. 4.3.3).
@@ 4.1 Experimental Setting
We use the implementations and the best hyperparameters provided by Yu et al. (2023) for all baseline models except CTAN. For CTAN, we use its official implementation, fixing the number of layers to 5. All models are trained with a batch size of 200 for fair efficiency analysis. For DyGMamba, we report the number of sampled one-hop temporal neighbors $\rho$ and the patch size $p$ here. On Wikipedia, Social Evo., and UCI, $\rho\ \&\ p=$ 32 & 1. On Reddit, $\rho\ \&\ p=$ 64 & 2. On MOOC, $\rho\ \&\ p$ = 128 & 4. On Enron, $\rho\ \&\ p=$256 & 8. On LastFM, $\rho\ \&\ p=$ 512 & 16. Note that to fairly compare DyGMamba’s efficiency with DyGFormer, we keep the sequence length $\rho/p$ input into the SSM as same as the length input into Transformer in Yu et al. (2023), i.e., $\rho/p=$ 32. All experiments are implemented with PyTorch (Paszke et al., 2019) on a server equipped with an AMD EPYC 7513 32-Core Processor and a single NVIDIA A40 with 45GB memory. We run each experiment for five times with five random seeds and report the mean results together with error bars. Further implementation details including complete hyperparameter configurations are presented in App. C. We employ two evaluation settings following previous works: the transductive and inductive settings. As suggested in (Poursafaei et al., 2022), we do link prediction evaluation using three negative sampling strategies (NSSs): random, historical and inductive. Historical NSS is only considered under the transductive setting. See App. D for detailed explanations. We employ two metrics, i.e., average precision (AP) and area under the receiver operating characteristic curve (AUC-ROC)
@@ 4.2 Performance Analysis
4.2.1 Comparative Study on Benchmark Datasets
@@ 4.2.2 Ablation Study
We conduct four ablation studies to study the effectiveness of model components. In study A, we make a model variant (Variant A) by removing the time-level SSM block and restrain our model from learning temporal patterns (information selection is substituted by mean pooling over the output of Eq. 4). In study B, we make a model variant (Variant B) by removing the Mamba SSM layers (Eq. 3) in the node-level SSM block. In study C, we switch the computation of the selection weights $\beta_{\theta}$ in Eq. 6d to $\beta_{\theta}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{\theta}))$ ($f_{\text{sel}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$) to create Variant C. In study D, we base on Variant C and develop Variant D that further enables information selection from opposite nodes, i.e., $\beta_{u}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{v}))$/ $\beta_{v}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{u}))$. Both ablation C and D do information selection without learning temporal patterns. From Table 3 and 4, we find that: (1) Variant A is constantly beaten by DyGMamba, showing the effectiveness of dynamic information selection based on edge-specific temporal patterns; (2) DyGMamba always outperforms Variant B, indicating the importance of encoding the one-hop temporal neighbors with SSM layers for capturing graph dynamics; (3) Variant C and D perform better than Variant A in most cases, implying that selecting temporal information is generally contributive; (4) Variant C generally lags behind Variant D, meaning that information selection from opposite node is beneficial; (5) DyGMamba performs better than both Variant C and D in almost all cases, proving that information selection based on temporal patterns is more effective.
@@
4.2.3 A Closer Look into Temporal Pattern Modeling with Synthetic Datasets
@@
We observe from ablation studies that dynamic information selection based on temporal patterns contributes to better model performance on real-world datasets. To better quantify its benefits, we construct three synthetic datasets, i.e., S1, S2 and S3, that follow different patterns and compare our model with DyGFormer, CTAN as well as Variant A, C, D on them. Each synthetic dataset contains 7 nodes, where the interactions of each pair of two nodes follow a certain pattern along time. And for each node, we generate interactions with all the other nodes. Assume we have a pair of node $u$ and $v$ and they have interactions at $\{t_{i}\}_{i=0}^{N}$, in S1, the time intervals between neighboring interactions $\{t_{1}-t_{0},...,t_{N}-t_{N-1}\}$ follow an increasing trend with a constant velocity of 0.05, i.e., $(t_{i+2}-t_{i+1})-(t_{i+1}-t_{i})=$ 0.05 . In S2, we set the time intervals to a decreasing trend with the same velocity, i.e., $(t_{i+1}-t_{i})-(t_{i+2}-t_{i+1})=$ 0.05. And in S3, we modify S1 by repeating several periods of increasing patterns taken from S1 to form a periodic dataset. In this way, we have three datasets demonstrating diverse temporal patterns: increasing/decreasing/periodic time intervals between neighboring interactions. Details of dataset construction and statistics are provided in App. A.2. From Table 5, we observe that DyGMamba greatly outperforms DyGFormer and CTAN. More importantly, Variant A, C and D show similar performance to DyGFormer, meaning that our time-level SSM block is able to capture temporal patterns and modeling such patterns for dynamic information selection is important in CTDG reasoning. For more implementation details on synthetic datasets, please refer to App. C.2.
@@
Datasets DyGFormer CTAN Variant A Variant C Variant D DyGMamba
S1 55.19 $\pm$ 0.98 51.25 $\pm$ 2.11 53.72 $\pm$ 0.04 55.45 $\pm$ 0.32 54.52 $\pm$ 0.71 81.58 $\pm$ 1.31
S2 57.80 $\pm$ 4.61 51.17 $\pm$ 0.93 60.16 $\pm$ 2.20 64.71 $\pm$ 2.33 61.51 $\pm$ 3.00 85.36 $\pm$ 2.55
S3 79.20 $\pm$ 0.60 51.46 $\pm$ 0.19 77.61 $\pm$ 2.31 77.61 $\pm$ 2.31 79.41 $\pm$ 2.13 86.59 $\pm$ 0.09
@@
Datasets DyGFormer CTAN Variant A Variant C Variant D DyGMamba
S1 56.27 $\pm$ 0.54 51.25 $\pm$ 2.38 53.16 $\pm$ 0.39 57.41 $\pm$ 0.04 55.83 $\pm$ 1.03 86.61 $\pm$ 1.30
S2 59.06 $\pm$ 6.07 51.46 $\pm$ 0.91 62.50 $\pm$ 2.28 64.93 $\pm$ 2.59 62.06 $\pm$ 3.40 89.94 $\pm$ 2.70
S3 82.89 $\pm$ 1.34 52.12 $\pm$ 0.44 81.78 $\pm$ 0.40 82.73 $\pm$ 1.97 84.20 $\pm$ 3.57 91.72 $\pm$ 0.11
@@ 4.3 Efficiency Analysis
We evaluate the models’ efficiency based on the following aspects: model size (number of trainable parameters), per-epoch training time, and GPU memory consumption during training. Since DyGMamba shares similarities with DyGFormer, i.e., both of them model large sequences of one-hop temporal neighbors and use patching to enhance scalability, we further analyze the impact of patch size on their scalability and performance. We also compare the complexity of DyGMamba and DyGFormer to highlight DyGMamba’s efficiency.
@@
4.3.1 Model Size, Per Epoch Training Time and GPU Memory Comparison Across Various Models
@@
Fig. 2 compares DyGMamba with five baselines in terms of number of parameters (model size), per epoch training time, and GPU memory consumption during training [^4] The baselines not included here are either extremely inefficient (e.g., CAWN) or inferior in performance (e.g., DyRep). Complete statistics of all baseline models presented in App. G. We find that: (1) DyGMamba uses very few parameters while maintaining the best performance, showing a strong parameter efficiency. Only CTAN constantly uses fewer parameters than DyGMamba, however, its performance is significantly worse with the only exception of Enron; (2) DyGMamba is always more efficient than DyGFormer with the same length of input sequence ($\rho/p=$32) while achieving the same performance; (3) Although DyGMamba generally consumes more GPU memory and takes more time to train per epoch compared with most baselines, the gap of consumption is modest. To model more temporal neighbors for long-range temporal dependent datasets, DyGMamba naturally requires more computational resources, thus enlarging the consumption gap. DyGFormer shows the same trend as DyGMamba since it also captures long-term temporal dependencies but at a higher cost; (4) CTAN requires very few computational resources. However, on long-range temporal dependent datasets, it is beaten by DyGFormer and DyGMamba by a large margin, e.g., on LastFM and Enron. Besides, CTAN is also hard to converge. Although it takes little time to train a single epoch, it needs more epochs to reach the best performance, leading to a long total training time. See App. G.2 for a total training time comparison among DyGFormer, CTAN and DyGMamba. To supplement, we also present in App. G.3 another experiment to compare DyGMamba with baselines in modeling an increasing number of temporal neighbors on Enron with limited total training time.
Figure 2: Efficiency comparison on four datasets among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory. The performance metric here is AP of transductive link prediction under random NSS. The greener, the better performance/efficiency. In contrast to other methods, DyGMamba consistently shows strong overall capability across different datasets. More explanations in Sec. 4.3.1.
@@
4.3.2 Impact of Patch Size on Scalability and Performance
Patching treats $p$ temporal neighbors as one patch and thus decreases the sequence length by $p$ times. This is very helpful in cutting the consumption of GPU memory and training/evaluation time. However, patching introduces excessive parameters because it is done through $f_{N}$, $f_{E}$, $f_{T}$ and $f_{F}$ whose sizes increase as the patch size grows. Fig. 3b shows the numbers of parameters of DyGFormer and DyGmamba with different patch sizes on a long-term temporal dependent dataset Enron. We find that patching greatly affects model sizes. To further study how patching affects DyGMamba, we decrease the patch size gradually from 8 to 1 and track DyGMamba’s performance (Fig. 3a) as well as efficiency (Fig. 3b to 3d) on Enron. Meanwhile, we also keep track on DyGFormer under the same patch size for comparison. We have several findings: (1) Whatever the patch size is, DyGMamba always consumes fewer parameters, less GPU memory and per epoch training time, showing its high efficiency; (2) While both models require increasing computational budgets as the patch size decreases, the speed of increase is much lower for DyGMamba, demonstrating its strong scalability in modeling longer sequences; (3) Different trends in performance change are observed between two models. While DyGFormer performs worse, DyGMamba can benefit from a smaller patch size, indicating its strong ability to capture nuanced temporal details even if the sequence becomes much longer. Note that the models use fewer parameters under smaller patch size. This also shows that DyGMamba can achieve much stronger parameter efficiency by reducing patch sizes. To further study the reason for finding (3), we plot the performance of Variant A under different patch sizes in Fig. 3a. We find that Variant A’s performance degrades when sequence length is more than $64$. This means that dynamic information selection based on edge-specific temporal patterns is essential for DyGMamba to optimally process long sequences. To supplement, we provide additional analysis on MOOC in App. J.
Reason for Diverse Performance Trends between DyGMamba and DyGFormer.
Patching decreases sequence length by applying linear transformation on a patch of node embeddings, giving a model much more parameters to tune (as indicated in Fig. 3b). Larger patch size mixes more sampled neighbors in each patch, introducing more training parameters while losing nuanced temporal details brought by the temporal order of the neighbors within each patch. For DyGFormer, the negative influence brought by mixing neighbors within patches is smaller than the positive influence brought by more trainable parameters, so the performance constantly increases with a growing patch size in Fig. 3a. By contrast, for DyGMamba, the negative influence brought by mixing neighbors is much greater than the positive influence brought by more trainable parameters, so it shows better performance when the patch size is smaller. For Variant A, given an increasing patch size, it follows the trend of DyGFormer when the patch size is below a threshold and shows degrading performance after that. This means that there is a trade-off between the lost temporal details and the additional parameters when we modify patch size. Also, by comparing Variant A and DyGMamba, we can tell that the difference in performance trend roots from the dynamic information selection module. Smaller patch size leads to longer sequences with more temporal details. The strong capability of the dynamic information selection module in long-range temporal reasoning let DyGMamba benefit from more nuanced temporal details, which is more influential than increasing training parameters.
4.3.3 Complexity Analysis: DyGMamba vs. DyGFormer
DyGMamba follows the current state-of-the-art DyGFormer by learning from one-hop temporal neighbors for temporal reasoning. We analyze the complexity of both models to show DyGMamba’s efficiency. Sequence length is the key factor affecting the consumption of computational resources in DyGFormer and DyGMamba. Following the computation of previous work (Zhu et al., 2024), the complexity of Transformer and Mamba in DyGFormer and DyGMamba can be written as
$O(\text{Transformer})=4(\rho/p)(4d)^{2}+2(\rho/p)^{2}(4d)=64(\rho/p)d^{2}+8d(\rho/p)^{2},$ (8a)
$O(\text{Mamba})=3(\rho/p)(4d)d_{\text{SSM}}+(\rho/p)(4d)d_{\text{SSM}}=16d_{\text{SSM}}d(\rho/p).$ (8b)
This means that DyGMamba holds a computational complexity linear to $\rho/p$, while DyGFormer’s complexity is quadratic to $\rho/p$. As a result, as the sequence length grows (either $\rho$ increases or $p$ decreases), DyGFormer is less scalable compared with DyGMamba [^5] As we set $k$ (number of edge-specific historical interactions discussed in Sec. 3.2) to a number much smaller than the number of sampled one-hop temporal neighbors, i.e., sequence length $\rho/p$, we omit here the contribution of the time-level SSM in complexity analysis.. Some may argue that increasing the patch size $p$ to a large enough value can offset the negative influence of the higher complexity of Transformer. However, as discussed in Sec. 4.3.2, increasing patch size will substantially increase model parameters, causing burden in parameter optimization, and meanwhile lose temporal details. This indicates that patching is not always beneficial and DyGMamba’s low complexity provides an alternative way to maintain great efficiency while considering more temporal information. With the same number of sampled temporal neighbors and equivalent computational resources, DyGMamba can leverage a smaller patch size, mitigating the negative effects of lost temporal details and simplifying parameter optimization, as implied in Sec. 4.3.2.
@@ 4.3.2 Impact of Patch Size on Scalability and Performance
Modeling Increasing Number of Temporal Neighbors with a Fixed Patch Size.
Sequence length is decided by the sampled temporal neighbors $\rho$ and the patch size $p$. If we want to model a huge number of temporal neighbors, e.g., $\rho=$ 4096, keeping the sequence length unchanged as 32 means we need to set $p$ to 128. Considering the potential negative impact brought by excessively large patch size (i.e., too many lost temporal details, harder optimization and much lower parameter efficiency), it is also important to see if a model is scalable to $\rho$, with a fixed $p$. In Fig. 4, we show the consumed GPU memory of DyGMamba and DyGFormer with $\rho$ varying from 64, 128, 256, 512, 1024, 2048, 4096 and 8192 on Enron, under a fixed patch size $p=$ 8. We find that DyGMamba shows superior scalability over DyGFormer. While DyGFormer can only process 2048 nodes, DyGMamba can deal with more than 8192 (at least 4 times) on a single 45GB NVIDIA A40 GPU. This implies our model’s potential to process a huge number of temporal neighbors with a limited GPU memory budget, while keeping a moderate patch size. The experiment provides a concrete example to demonstrate DyGMamba’s advantage in efficiency brought by its low complexity.